# Quaternionic Pole Placement via Companion Forms and the Ackermann Formula

### v1

@@ V Control Problem: State-Feedback Pole Placement in Quaternionic Systems
In the real or complex case, the method assigns $n$ (not necessarily distinct) eigenvalues of $A_{\mathrm{cl}}$. In the quaternionic setting, the situation is different: spectra must be interpreted through right eigenvalues (1) which are determined only up to quaternionic similarity (i.e., $\lambda\sim\alpha^{-1}\lambda\,\alpha$ for any $\alpha\in\mathbb{H}\setminus\{0\}$). Consequently, pole placement specifies the closed-loop right spectrum as a multiset of such similarity classes (each class being finite or infinite), or, equivalently, by choosing canonical complex representatives
$\lambda_{k}\in\mathbb{C},\quad\operatorname{Im}\lambda_{k}\geq 0,\quad k=1,\dots,n,$
counted with algebraic multiplicity. Under the usual controllability assumptions, the feedback gain $K$ can then be designed to realize the desired right-spectrum classes of $A_{\mathrm{cl}}$.
@@
Given $A\in\mathbb{H}^{n\times n}$, $B\in\mathbb{H}^{n}$, and desired targets $\{\lambda_{k}\}_{k=1}^{n}\subset\mathbb{H}$ (or their canonical complex representatives), find $K\in\mathbb{H}^{1\times n}$ such that, for the closed loop $A_{\rm cl}=A-BK$,
$\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$ $=$ $\textstyle{\bigcup}_{i}\,[\lambda_{i}],$
with similarity classes $\lambda_{k}]$ taken with the prescribed multiplicities.
Typically, the $\{\lambda_{k}\}$ are chosen as the right zeros of a desired monic closed-loop companion polynomial $a_{d}(\lambda)$. When $a_{d}\in\mathbb{R}[\lambda]$, nonreal roots appear in conjugate pairs and encode one spherical class each. For continuous time systems, pick representatives in the open left half–plane; $\operatorname{Re}$ sets decay rate and $\left|\operatorname{Im}\right|$ sets oscillation frequency.
@@ VI Quaternionic Pole Placement via Controllable Companion Form
In the real/complex setting, state feedback is obtained by transforming a controllable pair to companion form and matching coefficients. Over $\mathbb{H}$, the same workflow applies for single-input controllable systems: use the companion polynomial—read from the last row of the companion matrix and computed over $\mathbb{H}$—and interpret poles as right-eigenvalue similarity classes. Pole placement then reduces to a coefficient update in controller coordinates, followed by mapping the gain back through the similarity transformation. Crucially, this construction avoids determinants and characteristic/minimal polynomials—as well as Cayley–Hamilton and adjugates—which do not carry over cleanly to $\mathbb{H}$.
@@
Assume $(A,B)$ with $A\in\mathbb{H}^{n\times n}$, $B\in\mathbb{H}^{n}$ is controllable, and let $(A_{c},B_{c})$ be its (lower) controllable companion form via $T$ so that $A_{c}=T^{-1}AT$, $B_{c}=T^{-1}B$. Write the last row of $A_{c}$ as $\left[-a_{0},\dots,-a_{n-1}\right]$ and let $a(\lambda)=a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}$ be the companion polynomial. Given any monic
there exists a unique feedback u=Kxu=-Kx, $K\in\mathbb{H}^{1\times n}$, such that the right spectrum of $A_{\mathrm{cl}}:=A-BK$ equals the multiset of right zeros (similarity classes, with multiplicities) of $a_{d}$. Moreover
@@
Setting $k_{i}=d_{i}-a_{i}$ yields $a_{\mathrm{cl}}(\lambda)=a_{d}(\lambda)$. By the previously established spectral property of the companion form, the right zeros of $a_{\mathrm{cl}}$ are the right eigenvalues of $A_{\mathrm{cl},c}$, and therefore of $A_{\mathrm{cl}}$. Uniqueness follows since the coefficient update $k_{i}=d_{i}-a_{i}$ is bijective and $K\mapsto K_{c}=KT$ is bijective (invertible $T$). Finally $K=K_{c}T^{-1}$. ∎
@@
has the finite right spectrum $\sigma_{\mathscr{r}}(A_{\rm cl})=\{-1,-2\}$ as desired.
@@
$K_{c}$ $=$ $\begin{bmatrix}3+\mathbf{i}-\mathbf{j}+\mathbf{k}&3-\mathbf{i}+\mathbf{j}-\mathbf{k}\end{bmatrix}$,
@@
Here $\{-1\!+\!\mathbf{i},-1\!-\!\mathbf{i}\}\subset\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$; equivalently, the entire spherical (similarity) class $\left[-1+\mathbf{i}\right]$ is assigned as expected.
(c) Quaternionic coefficients: Demanding specific quaternion poles (e.g., $\lambda_{1}=-1+\mathbf{j}$ and $\lambda_{2}=-2+\mathbf{k}$) sets (rounded)
@@
Here $a_{d}(\lambda)\in\mathbb{H}[\lambda]$ has nonreal (noncentral) coefficients; evaluating $a_{d}(A)$ is then quaternionic, and we obtain
$K_{c}$ $=$ ${\begin{bmatrix}3.7\!-\!2\mathbf{i}\!-\!0.33\mathbf{j}\!-\!0.67\mathbf{k},&4\!+\!0.33\mathbf{i}\!-\!1.3\mathbf{j}\!+\!0.67\mathbf{k}\end{bmatrix}}$,
$K$ $=$ ${\begin{bmatrix}3.3\!+\!0.33\mathbf{j}\!+\!2.7\mathbf{k},&-2\!+\!1.7\mathbf{i}\!+\!0.33\mathbf{j}\!-\!0.67\mathbf{k}\end{bmatrix}}$,
@@
$A_{\rm cl}$ $=$ ${\begin{bmatrix}-2.3\!-\!0.33\mathbf{j}\!-\!2.7\mathbf{k},&2\!-\!0.67\mathbf{i}\!-\!0.33\mathbf{j}\!+\!0.67\mathbf{k}\\ 2.7\!+\!0.33\mathbf{i}\!+\!\mathbf{j}\!-\!3.3\mathbf{k},&\!-\!0.67\!+\!0.33\mathbf{i}\!-\!1.7\mathbf{j}\!+\!3\mathbf{k}\end{bmatrix}}$.
Its right spectrum is $\sigma_{\mathscr{r}}(A_{\mathrm{cl}})=\left[-1+\mathbf{i}\right]\cup[-2+\mathbf{i}]=\left[-1+\mathbf{j}\right]\cup\left[-2+\mathbf{k}\right]$ so $1\!+\!\mathbf{j},\,-2\!+\!\mathbf{k}\in\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$ confirming that the controller-form coefficient update achieves the prescribed classes even with quaternionic coefficients.
@@ VII Quaternionic Ackermann Formula
In this section, we show that the Ackermann formula carries over to quaternionic systems. Unlike the companion-form coefficient-update method, it is not fully general: it applies when the desired closed-loop companion polynomial has real coefficients. Practically, this means we assign right eigenvalue classes (spherical similarity classes) with chosen multiplicities—still broad enough for design, but not literally unrestricted. Algorithmically, the recipe is identical to the classical case (same controllability matrix selection and one-shot gain computation). At the same time, the proof is different in substance: over $\mathbb{H}$ we cannot appeal to the determinant or Cayley–Hamilton, so we justify the formula via the controllable companion relations and a central (real-coefficient) polynomial calculus.
@@
be the desired monic closed-loop companion polynomial with real coefficients and let $e_{n}=[0,\dots,0,1]^{\mathsf{T}}\in\mathbb{R}^{n}$ be the last canonical basis vector. Then the state-feedback row vector
@@
with u=Kxu=-Kx yields the closed-loop matrix ABKA-BK whose right spectrum equals the similarity classes induced by the right roots of $a_{d}(\lambda)$
$\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$ $=$ $\bigcup_{i=1}^{n}[\lambda_{i}],$ (35)
where $\{\lambda_{i}\}_{i=1}^{n}$ are the right roots of $a_{d}(\lambda)$ (listed with multiplicity).
@@
For an $n$th-order single-input design, each real class uses one degree and each nonreal class uses two; with $r$ real and $s$ nonreal classes, $2s+r\leq n.$
@@
need not hold once any coefficient of $p$ is a nonreal quaternion. Indeed, take, $A=\mathbf{i}$ and $T=\mathbf{k}$ so that $A_{c}=T^{-1}AT=-\mathbf{i},$ and $p(\lambda)=\lambda+\mathbf{i}$. We have $p(A)T=(\mathbf{i}+\mathbf{i})k=2\mathbf{i}\mathbf{k}=-2\mathbf{j}$ but $Tp(A_{c})=k(-\mathbf{i}+\mathbf{i})=0$, hence $p(A)T\neq Tp(A_{c})$. This illustrates why the Ackermann proof below requires $p\in\mathbb{R}[\lambda]$: real coefficients ensure unambiguous evaluation and permit the coefficient–matrix “swap” used in the intertwining step.
@@
Classical proofs over $\mathbb{R}$ or $\mathbb{C}$ lean on determinants and adjoints [11, 4], characteristic polynomials [4, 8], or Cayley–Hamilton—steps [11, 4]; see also Ackermann’s original paper [1]. These steps do not transfer verbatim to $\mathbb{H}$ because of noncommutativity and left/right evaluation. We therefore avoid those tools and proceed differently:
@@
Given the desired (real) closed-loop polynomial
@@
assigns the right eigenvalues of $A_{c}-B_{c}K_{c}$ to the roots of $a_{d}$ (counted with multiplicity, interpreted as right-eigenvalue classes). The feedback in the original coordinates is
$K=K_{c}T^{-1}$
Step 2: Controllable-form Ackermann equals coefficient matching. Let $\mathcal{C}_{c}=\left[B_{c},A_{c}B_{c},\ldots,A_{c}^{n-1}B_{c}\right]$ be the controllability matrix of the controllable form (4), and $e_{n}$ the last canonical basis vector. The controllable-form Ackermann expression
@@
produces exactly the same $K_{c}$ as above. The special structure of $\mathcal{C}_{c}^{-1}$ in the controllable companion basis implies $e_{n}^{\top}\mathcal{C}_{c}^{-1}=e_{1}^{\top}$. Using this in (42), we have
@@
Furthermore, the shape of $A_{c}$ gives
$e_{1}^{\mathsf{T}}A_{c}^{k}$ $=$ $e_{k+1}^{\top}\quad(k=0,\dots,n-2),$
$e_{1}^{\mathsf{T}}A_{c}^{n}$ $=$ $\big[-a_{0},\dots,-a_{n-1}\big],$
so, since $e_{1}$ commutes with the $d_{i}$ (all real),
@@
$=$ $\begin{bmatrix}d_{0}-a_{0}&\ldots&d_{n-1}-a_{n-1}\end{bmatrix}$
@@
Step 3 — Coordinate-free form and mapping back: Let $\mathcal{C}$ be the controllability matrix (3) of $(A,B)$. From the similarity, for $k\geq 0$ we have
$B$ $=$ $TB_{c},\quad A^{k}T=TA_{c}^{\,k}.$ (43)
@@
$\mathcal{C}=T\mathcal{C}_{c}\implies\mathcal{C}^{-1}=\mathcal{C}_{c}^{-1}T^{-1}\;\mathrm{and\;}\;T=\mathcal{C}\,\mathcal{C}_{c}^{-1}.$ (44)
Because $a_{d}\in\mathbb{R}[\lambda],$ the Lemma 1 yields
$T^{-1}\,a_{d}(A)$ $=$ $a_{d}(A_{c})\,T^{-1}.$ (45)
@@
$K$ $=$ $e_{n}^{\mathsf{T}}\mathcal{C}^{-1}a_{d}(A)=e_{n}^{\top}C_{c}^{-1}\,T^{-1}a_{d}(A)$
$=$ $\big(e_{n}^{\top}C_{c}^{-1}a_{d}(A_{c})\big)\,T^{-1}=K_{c}T^{-1}.$
since $K_{c}=e_{n}^{\mathsf{T}}C_{c}^{-1}a_{d}(A_{c}).$
Thus, the basis-invariant identity (VII) produces exactly the same feedback as the controllable-form construction, completing the proof. ∎
@@
which equals (28). The resulting closed-loop matrix $A_{\rm cl}=A-BK$ is identical to (29) and its right spectrum is $\sigma_{\mathscr{r}}(A_{\rm cl})=\{-1,-2\}$ as desired.
@@
which equals (30). The closed-loop matrix $A_{\rm cl}$ is identical to (31) and $\{-1\!+\!\mathbf{i},-1\!-\!\mathbf{i}\}\subset\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$.
(c) Quaternionic coefficients, for which the formula fails: Demanding quaternion poles $\lambda_{1}=-1+\mathbf{j}$, $\lambda_{2}=-2+\mathbf{k}$ and using $a_{d}(\lambda)$ as in (32), the Ackermann returns
$K$ $=$ ${\begin{bmatrix}3.5\!+\!1.3\mathbf{i}\!-\!0.5\mathbf{j}\!+\!3\mathbf{k},&-1.5\!+\!2\mathbf{i}\!-\!0.5\mathbf{j}\!-\!1.7\mathbf{k}\end{bmatrix}}$.
@@
$A_{\mathrm{cl}}$ $=$ ${\begin{bmatrix}-2.5\!-\!1.3\mathbf{i}\!+\!0.5\mathbf{j}\!-\!3\mathbf{k},&1.5\!-\!\mathbf{i}\!+\!0.5\mathbf{j}\!+\!1.7\mathbf{k}\\[2.0pt] 3\!-\!0.5\mathbf{i}\!-\!0.33\mathbf{j}\!-\!3.5\mathbf{k},&1.7\!-\!0.5\mathbf{i}\!-\!2\mathbf{j}\!+\!2.5\mathbf{k}\end{bmatrix}}$
whose right spectrum is $0.48+0.85\mathbf{i}\,]\cup[\,-3.7+2.4\mathbf{i}\,]$, confirming that when $a_{d}\notin\mathbb{R}[\lambda]$, the Ackermann construction does not enforce the prescribed classes.

### v2

@@ V Control Problem: State-Feedback Pole Placement in Quaternionic Systems
In the real or complex case, the method assigns $n$ (not necessarily distinct) eigenvalues of $A_{\mathrm{cl}}$. In the quaternionic setting, the situation is different: spectra must be interpreted through right eigenvalues (1) which are determined only up to quaternionic similarity (i.e., $\lambda\sim\alpha^{-1}\lambda\,\alpha$ for any $\alpha\in\mathbb{H}\setminus\{0\}$). Consequently, pole placement specifies the closed-loop right spectrum as the set of such similarity classes. Equivalently, one may specify canonical complex representatives
$\lambda_{k}\in\mathbb{C},\qquad\operatorname{Im}\lambda_{k}\geq 0,\qquad k=1,\dots,n,$
counted with algebraic multiplicity.
@@
Given $A\in\mathbb{H}^{n\times n}$, $B\in\mathbb{H}^{n}$, and desired representatives $\lambda_{1},\dots,\lambda_{n}\in\mathbb{H}$ (not necessarily distinct), or equivalently their canonical complex representatives, find $K\in\mathbb{H}^{1\times n}$ such that, for the closed loop $A_{\rm cl}=A-BK$,
$\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$ $=$ $\{[\mu_{1}],\dots,[\mu_{m}]\},$
where $\mu_{1},\dots,\mu_{m}$ are representatives of the distinct desired similarity classes appearing among the targets. Any desired multiplicities are encoded separately by the chosen polynomial data.
Typically, one chooses a monic closed-loop companion polynomial $a_{d}(\lambda)$ whose right zeros represent the desired similarity classes. Thus $a_{d}$ is a convenient algebraic encoding of the pole-placement target, while the actual control objective is the closed-loop right spectrum. When $a_{d}\in\mathbb{R}[\lambda]$, nonreal zeros appear in conjugate pairs and encode one spherical class each. For continuous-time systems, one usually picks canonical complex representatives in the open left half-plane; $\operatorname{Re}$ sets the decay rate and $|\operatorname{Im}|$ the oscillation frequency.
@@ VI Quaternionic Pole Placement via Controllable Companion Form
In the real/complex setting, state feedback is obtained by transforming a controllable pair to companion form and matching coefficients. Over $\mathbb{H}$, the same workflow applies for single-input controllable systems: use the companion polynomial—read from the last row of the companion matrix and computed over $\mathbb{H}$—and interpret poles as right-eigenvalue similarity classes. Pole placement then reduces to a coefficient update in controllable-companion coordinates, followed by mapping the gain back through the similarity transformation. Crucially, this construction avoids determinants and characteristic/minimal polynomials—as well as Cayley–Hamilton and adjugates—which do not carry over cleanly to $\mathbb{H}$.
@@
Assume $(A,B)$ with $A\in\mathbb{H}^{n\times n}$, $B\in\mathbb{H}^{n}$ is controllable, and let $(A_{c},B_{c})$ be its (lower) controllable companion form via $T$ so that $A_{c}=T^{-1}AT$, $B_{c}=T^{-1}B$. Write the last row of $A_{c}$ as $\left[-a_{0},\dots,-a_{n-1}\right]$ and let $a(\lambda)=a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}$ be the companion polynomial. Given any monic quaternionic polynomial
$a_{d}(\lambda)$ $=$ $d_{0}+d_{1}\lambda+\cdots+d_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{H}[\lambda],$ (19)
there exists a unique feedback u=Kxu=-Kx, $K\in\mathbb{H}^{1\times n}$, such that the closed-loop companion polynomial of $A_{\mathrm{cl},c}:=A_{c}-B_{c}K_{c}$, with $K_{c}:=KT$, equals $a_{d}$. Consequently, the right spectrum of $A_{\mathrm{cl}}:=A-BK$ is the set of similarity classes determined by the right zeros of $a_{d}$. The multiplicities are encoded by the polynomial $a_{d}$ itself, not by the set-valued spectrum. Moreover
@@
Setting $k_{i}=d_{i}-a_{i}$ for $i=0,\dots,n-1$ yields $a_{\mathrm{cl}}(\lambda)=a_{d}(\lambda)$. By the previously established spectral property of the companion form, the right spectrum of $A_{\mathrm{cl},c}$ is the set of similarity classes determined by the right zeros of $a_{\mathrm{cl}}$, hence by those of $a_{d}$. Since $A_{\mathrm{cl}}$ is similar to $A_{\mathrm{cl},c}$, the same holds for $A_{\mathrm{cl}}$.
Uniqueness holds for the prescribed polynomial $a_{d}$: the coefficient update $k_{i}=d_{i}-a_{i}$ is bijective, and $K\mapsto K_{c}=KT$ is bijective because $T$ is invertible. Hence there is a unique feedback whose closed-loop companion polynomial equals $a_{d}$. Finally $K=K_{c}T^{-1}$. ∎
@@
has the right spectrum $\sigma_{\mathscr{r}}(A_{\rm cl})=\{[-1],[-2]\}$ as desired.
@@
$K_{c}$ $=$ $\begin{bmatrix}3-\mathbf{i}+\mathbf{j}-\mathbf{k}&3+\mathbf{i}-\mathbf{j}+\mathbf{k}\end{bmatrix}$,
@@
Here $\sigma_{\mathscr{r}}(A_{\mathrm{cl}})=\{[-1+\mathbf{i}]\}$; equivalently, $1\pm\mathbf{i}$ are the two canonical complex representatives of the assigned spherical class.
(c) Quaternionic coefficients: All displayed numerical values in this example are rounded independently from full-precision computations; therefore, the printed intermediate quantities are not intended to reproduce one another exactly. Demanding quaternionic target classes, represented for example by $\lambda_{1}=-1+\mathbf{j}$ and $\lambda_{2}=-2+\mathbf{k}$, yields the following rounded target polynomial
@@
Here $a_{d}(\lambda)\in\mathbb{H}[\lambda]$ has nonreal (noncentral) coefficients, but the companion-form coefficient update still applies. Using $K_{c}=[d_{0}-a_{0},\dots,d_{n-1}-a_{n-1}]$, we obtain
$K_{c}$ $=$ ${\begin{bmatrix}3.7\!-\!2\mathbf{i}\!-\!0.33\mathbf{j}\!-\!0.67\mathbf{k}&4\!+\!0.33\mathbf{i}\!-\!1.3\mathbf{j}\!+\!0.67\mathbf{k}\end{bmatrix}}$,
$K$ $=$ ${\begin{bmatrix}3.3\!+\!0.33\mathbf{j}\!+\!2.7\mathbf{k}&-2\!+\!1.7\mathbf{i}\!+\!0.33\mathbf{j}\!-\!0.67\mathbf{k}\end{bmatrix}}$,
@@
$A_{\rm cl}$ $=$ ${\begin{bmatrix}-2.3\!-\!0.33\mathbf{j}\!-\!2.7\mathbf{k}&2\!-\!0.67\mathbf{i}\!-\!0.33\mathbf{j}\!+\!0.67\mathbf{k}\\ 2.7\!+\!0.33\mathbf{i}\!+\!\mathbf{j}\!-\!3.3\mathbf{k}&\!-\!0.67\!+\!0.33\mathbf{i}\!-\!1.7\mathbf{j}\!+\!3\mathbf{k}\end{bmatrix}}$.
Its right spectrum is
$\sigma_{\mathscr{r}}(A_{\mathrm{cl}})=\{[-1+\mathbf{i}],\,[-2+\mathbf{i}]\}=\{[-1+\mathbf{j}],\,[-2+\mathbf{k}]\}.$
confirming that the controllable-companion-form coefficient update achieves the prescribed classes even with quaternionic coefficients.
@@ VII Quaternionic Ackermann Formula
In this section we derive a quaternionic Ackermann formula. Unlike the companion-form coefficient update, it applies only when the chosen target polynomial $a_{d}$ has real (central) coefficients. This still covers real classes and spherical classes arising from complex-conjugate pairs, while the proof proceeds via controllable companion relations rather than determinants or Cayley–Hamilton.
Remark 6.
Throughout this section, the desired polynomial $a_{d}$ is assumed to belong to $\mathbb{R}[\lambda]$. Since its coefficients are real and hence central, the left and right scalar evaluations coincide, and the same holds for the corresponding matrix evaluations. Accordingly, we write simply $a_{d}(q)$ and $a_{d}(A)$, without the subscripts $\ell$ or $r$. This restriction is essential for the Ackermann formula below; for quaternionic coefficients, the formula need not hold.
@@
be a chosen monic real polynomial whose zeros encode the desired closed-loop similarity classes and let $e_{n}=[0,\dots,0,1]^{\mathsf{T}}\in\mathbb{R}^{n}$ be the last canonical basis vector. Then the state-feedback row vector
@@
with u=Kxu=-Kx yields the closed-loop matrix ABKA-BK whose right spectrum is the set of similarity classes determined by the zeros of $a_{d}(\lambda)$:
$\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$ $=$ $\{[\lambda]:a_{d}(\lambda)=0\}.$ (29)
Equivalently, if $\lambda_{1},\dots,\lambda_{m}$ are representatives of the distinct similarity classes of zeros of $a_{d}$, then
$\sigma_{\mathscr{r}}(A_{\mathrm{cl}})=\{[\lambda_{1}],\dots,[\lambda_{m}]\}.$
The algebraic multiplicities are encoded by the polynomial $a_{d}$ itself, not by the set $\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$.
@@
For an $n$th-order single-input design, each real zero contributes one degree, while each nonreal conjugate pair contributes two. Counting algebraic multiplicity, these contributions sum to $n$.
@@
need not hold once any coefficient of $p$ is a nonreal quaternion. Indeed, take $A=\mathbf{i}$ and $T=\mathbf{k}$ so that $A_{c}=T^{-1}AT=-\mathbf{i},$ and $p(\lambda)=\lambda+\mathbf{i}$. We have $p(A)T=(\mathbf{i}+\mathbf{i})k=2\mathbf{i}\mathbf{k}=-2\mathbf{j}$ but $Tp(A_{c})=k(-\mathbf{i}+\mathbf{i})=0$, hence $p(A)T\neq Tp(A_{c})$. This illustrates why the Ackermann proof below requires $p\in\mathbb{R}[\lambda]$: real coefficients ensure unambiguous evaluation and permit the coefficient–matrix “swap” used in the intertwining step.
@@
Classical proofs over $\mathbb{R}$ or $\mathbb{C}$ lean on determinants and adjoints [10, 4], characteristic polynomials [4, 7], or Cayley–Hamilton; see also Ackermann’s original paper [1]. These arguments do not transfer verbatim to $\mathbb{H}$ because of noncommutativity and left/right evaluation. We therefore proceed differently.
@@
Given the desired real monic closed-loop polynomial
@@
assigns, by Theorem 3, the right spectrum of $A_{c}-B_{c}K_{c}$ to the set of similarity classes determined by the right roots of $a_{d}$. The multiplicities are encoded in the factorization of $a_{d}$, not in the set-valued spectrum $\sigma_{\mathscr{r}}(\cdot)$. The feedback in the original coordinates is
$K$ $=$ $K_{c}T^{-1}.$
Step 2: Controllable-form Ackermann equals coefficient matching. Let $\mathcal{C}_{c}=\left[B_{c},A_{c}B_{c},\ldots,A_{c}^{n-1}B_{c}\right]$ be the controllability matrix of the controllable form (4). The controllable-form Ackermann expression
@@
produces exactly the same $K_{c}$ as above. Indeed, applying Theorem 1 to the already-companion pair $(A_{c},B_{c})$ gives $T=I$, hence
$e_{n}^{\top}\mathcal{C}_{c}^{-1}=e_{1}^{\top}.$
Therefore
@@
Furthermore, the shape of $A_{c}$ yields
$e_{1}^{\mathsf{T}}A_{c}^{k}$ $=$ $e_{k+1}^{\top},\qquad k=0,\dots,n-1,$
$e_{1}^{\mathsf{T}}A_{c}^{n}$ $=$ $e_{1}^{\mathsf{T}}A_{c}^{n-1}A_{c}=e_{n}^{\top}A_{c}=\big[-a_{0},\dots,-a_{n-1}\big].$
Since the coefficients $d_{i}$ are real and hence central,
@@
$=$ $\begin{bmatrix}d_{0}-a_{0}&\ldots&d_{n-1}-a_{n-1}\end{bmatrix}$,
@@
Step 3: Coordinate-free form and mapping back. Let $\mathcal{C}$ be the controllability matrix (3) of $(A,B)$. From the similarity, for $k\geq 0$ we have
$B$ $=$ $TB_{c},\qquad A^{k}T=TA_{c}^{\,k}.$
@@
$\mathcal{C}=T\mathcal{C}_{c}\;\Longrightarrow\;\mathcal{C}^{-1}=\mathcal{C}_{c}^{-1}T^{-1},\qquad T=\mathcal{C}\mathcal{C}_{c}^{-1}.$ (37)
Because $a_{d}\in\mathbb{R}[\lambda]$, Lemma 1 gives
$T^{-1}a_{d}(A)$ $=$ $a_{d}(A_{c})T^{-1}.$
@@
$K$ $=$ $e_{n}^{\mathsf{T}}\mathcal{C}^{-1}a_{d}(A)=e_{n}^{\top}\mathcal{C}_{c}^{-1}T^{-1}a_{d}(A)$ (38)
$=$ $\big(e_{n}^{\top}\mathcal{C}_{c}^{-1}a_{d}(A_{c})\big)T^{-1}=K_{c}T^{-1},$
since $K_{c}=e_{n}^{\mathsf{T}}\mathcal{C}_{c}^{-1}a_{d}(A_{c})$.
Thus, the basis-invariant formula (28) produces exactly the same feedback as the controllable-form construction, completing the proof. ∎
@@
which equals (22). The resulting closed-loop matrix $A_{\rm cl}=A-BK$ is identical to (23) and its right spectrum is $\sigma_{\mathscr{r}}(A_{\rm cl})=\{[-1],[-2]\}$ as desired.
@@
which equals (24). The closed-loop matrix $A_{\rm cl}$ is identical to (25) and $\sigma_{\mathscr{r}}(A_{\mathrm{cl}})=\{[-1+\mathbf{i}]\}$.
(c) Quaternionic coefficients, for which the formula fails: All displayed numerical values in this example are rounded independently from full-precision computations; therefore, the printed intermediate quantities are not intended to reproduce one another exactly. The Ackermann formula cannot assign quaternionic target classes represented, for example, by $\lambda_{1}=-1+\mathbf{j}$ and $\lambda_{2}=-2+\mathbf{k}$. If one nevertheless applies (28), incorrectly interpreting
$a_{d}(A)\quad\text{as}\quad a_{d}(A)_{\mathscr{r}}=d_{0}I+d_{1}A+A^{2},$
then Ackermann returns
$K$ $=$ ${\begin{bmatrix}3.5\!+\!1.3\mathbf{i}\!-\!0.5\mathbf{j}\!+\!3\mathbf{k}&-1.5\!+\!2\mathbf{i}\!-\!0.5\mathbf{j}\!-\!1.7\mathbf{k}\end{bmatrix}}$.
@@
$A_{\mathrm{cl}}$ $=$ ${\begin{bmatrix}-2.5\!-\!1.3\mathbf{i}\!+\!0.5\mathbf{j}\!-\!3\mathbf{k}&1.5\!-\!\mathbf{i}\!+\!0.5\mathbf{j}\!+\!1.7\mathbf{k}\\[2.0pt] 3\!-\!0.5\mathbf{i}\!-\!0.33\mathbf{j}\!-\!3.5\mathbf{k}&-1.7\!-\!0.5\mathbf{i}\!-\!2\mathbf{j}\!+\!2.5\mathbf{k}\end{bmatrix}}$
whose right spectrum is $\{[-0.48+0.85\mathbf{i}],\,[-3.7+2.4\mathbf{i}]\}$, confirming that when $a_{d}\notin\mathbb{R}[\lambda]$, the Ackermann construction does not enforce the prescribed classes.